OpenStudy (anonymous):

Identify the vertex for the graph of y = −3x2 + 6x − 2. (1, 1) (−1, 1) (1, −5) (−1, −5)

OpenStudy (anonymous):

\[-\frac{b}{2a}=-\frac{6}{2\times (-3)}=1\]

OpenStudy (anonymous):

so first coordinate is 1, second coordinate is what you get when you replace \(x\) by \(1\)